OpenStudy (anonymous):

help me plz fill in the blank with the numerator of the difference shown below. x/x+2 - 1/x+2 =??/x+2

OpenStudy (anonymous):

\[\frac{x}{x+2}-\frac{1}{x+2}\] right ?

OpenStudy (anonymous):

since the denominators are the same, you can subtract straight across and get \[\frac{x-1}{x+2}\]

OpenStudy (anonymous):

oh got it !! thank you :)